# Flat-band FFLO State from Quantum Geometric Discrepancy

### v2

@@ Abstract
We propose a new scheme for the Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) Cooper pairing states within ultraflat bands. Central to our approach is the concept of “quantum geometric discrepancy” (QGD), which characterizes the discrepancy in the quantum geometry of paired electrons, giving rise to the flat-band FFLO instability. Remarkably, we find that this instability is directly related to a new quantum geometric quantity we term “anomalous quantum distance”, which formally captures QGD. To model both QGD and the anomalous quantum distance, we examine a flat-band electronic Hamiltonian with tunable spin-dependent quantum metrics. Utilizing the band-projection method, we analyze the QGD-induced FFLO instability from pair susceptibility near the superconducting critical temperature. Furthermore, we perform self-consistent mean-field calculations to obtain the phase diagram of the BCS-FFLO transition driven by QGD, which aligns well with our analytical results. We emphasize that QGD serves as a distinctive protocol for stabilizing the flat-band FFLO phase.
@@
Introduction.—Understanding unconventional superconductivity is a central theme in modern condensed matter physics, with important implications for both theoretical and experimental aspects Sigrist and Ueda (1991); Mineev et al. (1999); Bauer and Sigrist (2012); Scalapino (2012). Among the various mechanisms that extend beyond the conventional Bardeen-Cooper-Schrieffer (BCS) theory of pairing Bardeen et al. (1957), the Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) phase is notable for its unique characterization of Cooper pairs with finite center-of-mass momentum $\bm{Q}$ Fulde and Ferrell (1964); Larkin and Ovchinnikov (1965); Kinnunen et al. (2018); Agterberg et al. (2020). Various mechanisms have been explored to realize these exotic FFLO states in systems characterized by dispersive bands Lu et al. (2015); Xi et al. (2016); Zhou et al. (2016); Saito et al. (2016); Liu (2017); Song and Koshelev (2019); Xie and Law (2023); Yuan (2023); Nakamura et al. (2024); Wan et al. (2023).
@@
In the original formulation of the FFLO phase, a spin-singlet pairing state with finite center-of-mass momentum $\bm{Q}$ ($|\bm{Q}|\propto B/v_{F}$) is favored in the presence of a large Zeeman field $B$ [see Fig. 1(a)]. However, the applicability of this microscopic mechanism becomes problematic in flat-band systems, where the Fermi velocity $v_{F}$ is nearly zero. This results in an unphysical divergence of $\bm{Q}$ as $v_{F}\rightarrow 0$, in the presence of a finite magnetic field. Consequently, two critical questions arise: First, what mechanism can stabilize the FFLO state as $v_{F}\rightarrow 0$? Second, what determines the center-of-mass momentum $\bm{Q}$ of Cooper pairs in the flat-band FFLO state? In dispersive bands, this is typically governed by Fermi surface nesting; however, in flat bands, the poorly defined Fermi surface weakens this relationship.
@@
Figure 1: Schematic illustrations of (a) conventional FFLO from Zeeman effect, and (b) flat-band FFLO from QGD. In (a), an external $B$ field induces spin-split Fermi surfaces, leading to the formation of Cooper pairs with finite momentum $\bm{Q}$, where $|\bm{Q}|\propto B/v_{F}$. In (b), the finite momentum $\bm{Q}$ is stabilized when the quantum metrics of the paired electrons are different ($\mathcal{G}_{\uparrow}\neq\mathcal{G}_{\downarrow}$) within flat bands.
@@
Lying in the context of the recent studies of the pronounced quantum geometric effects in flat-band systems Peotta and Törmä (2015); Julku et al. (2016); Liang et al. (2017); Hu et al. (2019); Julku et al. (2020); Xie et al. (2020); Peri et al. (2021); He et al. (2021); Herzog-Arbeitman et al. (2022); Törmä et al. (2022, 2018); Huhtinen et al. (2022); Hofmann et al. (2023); Jiang and Barlas (2023); Chen and Huang (2023); Han et al. (2024); Ticea et al. (2024); Wang and Huang (2024); Chen and Law (2024); Hu et al. (2023); Kitamura et al. (2022); Törmä (2023); Tian et al. (2023); Yu et al. (2024), in this work, we address the aforementioned questions by linking the FFLO instability to the quantum geometry of Bloch wave functions. For concreteness, we propose a new scheme for FFLO pairing states within an isolated flat band, driven by “quantum geometric discrepancy (QGD)”. This concept refers to the discrepancy in the quantum geometry of the wave functions of the two electrons forming a Cooper pair, which explicitly breaks time reversal ($\mathcal{T}$) symmetry. Based on the band-projection formalism, we analyse the static pair susceptibility and establish that the critical temperature for finite-momentum pairing, which leads to FFLO instability, is directly associated with the band average of a quantum geometric quantity [as described in Eq. (6)]—“anomalous quantum distance (AQD)” $\bar{d}_{\bm{k}^{\prime},\bm{k}}$, formally defined as
@@
Here, $u_{\bm{k}\sigma}$ represents the Bloch wave function at momentum $\bm{k}$ and spin $\sigma$. AQD quantifies the overlap between the wave functions of paired electrons. In the presence of $\mathcal{T}$ symmetry, where $u_{\bm{k}\uparrow}=u^{*}_{-\bm{k}\downarrow}$, the quantity $\bar{d}_{\bm{k},\bm{k}^{\prime}}$ becomes the well-known Hilbert–Schmidt quantum distance Shapere and Wilczek (1989); Bužek and Hillery (1996); Dodonov et al. (2000); Rhim et al. (2020), which measures the quantum mechanical distance between two Bloch states. In this scenario, the BCS (zero-momentum) pairing state is favored. In the absence of $\mathcal{T}$ symmetry, $u_{\bm{k}\uparrow}\neq u^{*}_{-\bm{k}\downarrow}$ generally makes this relation unestablished, and Eq. (1) will exhibit anomalous behaviors due to the QGD, enabling a finite-momentum pairing state.
As we will demonstrate in the following sections, by examining a flat-band model with tunable spin-dependent quantum metrics, we find that even in the flat-band limit, the finite momentum $\bm{Q}$ can be stabilized when pairing is frustrated by QGD, specifically when $\mathcal{G}_{\uparrow}\neq\mathcal{G}_{\downarrow}$ [see Fig. 1(b)], where $\mathcal{G}_{\uparrow}$ and $\mathcal{G}_{\downarrow}$ denote the quantum metrics of the Bloch functions of the paired electrons. Our findings provide a viable route toward realizing flat-band FFLO states.
Effective action from band projection.—To investigate the FFLO instability on the flat band, we first derive the static superconducting Gaussian fluctuations using the band-projection formalism Setty et al. (2023); Iskin (2023); Chen and Law (2024), and this analysis is rather general when considering conventional pairing. Without loss of generality, we consider the local Hubbard interaction $H_{\mathrm{int}}=-U\sum_{i\alpha}\hat{c}^{\dagger}_{i\alpha\uparrow}\hat{c}^{\dagger}_{i\alpha\downarrow}\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}$, where $U>0$, $\alpha$ and $i$ denote the orbital and site index.
In order to analyze the low-energy behavior of superconducting pairings on the relevant isolated band using field theory, after the Fourier transformation ($\hat{c}_{i\alpha\sigma}=1/\sqrt{N_{c}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{r}_{i}}\hat{c}_{\alpha,\bm{k}\sigma}$, where $N_{c}$ is the number of unit cells) and the uniform pairing approximation Törmä et al. (2018); Huhtinen et al. (2022), we project the interaction onto the target flat band by $\hat{c}_{\alpha,\bm{k}\sigma}\rightarrow u_{\bm{k}\sigma}^{*}(\alpha)\hat{a}_{\bm{k}\sigma}$, where $\hat{c}_{\alpha,\bm{k}\sigma}$ and $\hat{a}_{\bm{k}\sigma}$ denotes the orbital and band-basis electron annihilation operators. We can obtain the effective attractive interaction
@@
in which $\Lambda(\bm{k},\bm{q})=\sum_{\alpha}u_{\bm{k}+\bm{q}\uparrow}(\alpha)u_{-\bm{k}\downarrow}(\alpha)$ is the form factor, which encodes the information of quantum geometry Chen and Law (2024); Hu et al. (2023). Here the effective coupling constant $g\approx U/N_{\mathrm{orb}}N_{c}$, where $N_{\mathrm{orb}}$ is the number of orbitals in each unit cell. For convenience, we naturally introduce the ideal condition such that the attractive interaction strength $g$ is large compared to the bandwidth $W$ of the isolated flat band but small compared to the band gap $\Delta_{\mathrm{gap}}$ ($W\ll g\ll\Delta_{\mathrm{gap}}$). From Eq. (2), we can arrive at an effective one-band description in the functional integral formalism Altland and Simons (2023), and the effective action can be expressed as
@@
Here we have used the notations $k=\left(\bm{k},\omega_{n}\right)$ and $q=\left(\bm{q},\Omega_{m}\right)$ with $\omega_{n}$ $(\Omega_{m})$ being the fermionic (bosonic) Matsubara frequencies. Then we adopt the standard Hubbard-Stratonovich decoupling of Eq. (3) to derive the effective action $S[\Delta_{q}]$ (for details see Supplemental Material Note I sup ) of the superconducting order parameter field $\Delta_{q}$. By retaining the second order terms of $\Delta_{q}$ in $S[\Delta_{q}]$, the Gaussian action around the normal state saddle point $\Delta_{q}=0$ reads
$S_{G}\left[\Delta_{q}\right]=\sum_{q}\Gamma_{q}^{-1}|\Delta_{q}|^{2},$ (4)
where the coefficient $\Gamma_{q}^{-1}=g^{-1}-\chi^{c}_{q}$ is the pairing propagator, and the static pair susceptibility $\chi^{c}_{\bm{q}}\equiv\chi^{c}_{q=(\bm{q},0)}$ reads
@@
Here we use the notation as $\xi_{\bm{k}}=\xi_{\bm{k}\uparrow}=\xi_{-\bm{k}\downarrow}$ and $n_{\mathrm{F}}$ is the Fermi-Dirac distribution. The number equation near the critical temperature is $N=-\beta^{-1}\partial S_{G}[\Delta_{q}=0]/\partial\mu$, where $N$ is the total particle number. This yields $2\mathcal{A}\nu=\sum_{\bm{k}}\left[1-\tanh\left(\beta\xi_{\bm{k}}/2\right)\right]$, where $\mathcal{A}$ is the area of the first Brillouin zone, $\beta^{-1}=k_{B}T$, and $\nu$ is the filling factor of the target band ($0\leq\nu\leq 1$).
Critical temperature of FFLO instability.—By solving the linearized gap equation $g\chi^{c}_{\bm{q}}=1$ together with the number equation, we can derive the critical temperature $T_{c\bm{q}}$ for the instability of a static, spatially non-uniform pairing state characterized by a finite $\bm{q}$ index. Restricting the analysis to single-$\bm{q}$ states, $\Delta_{\bm{q}}\propto\delta_{\bm{q},\bm{Q}}$, the value of $\bm{Q}$ is found at the highest $T_{c\bm{q}}$. If $\bm{Q}\neq 0$ ($\bm{Q}=0$), the pairing instability is associated with the FFLO (BCS) phase.
Importantly, in the flat-band limit, we find $\chi^{c}_{\bm{q}}=\beta\sum_{\bm{k}}|\Lambda(\bm{k},\bm{q})|^{2}(1/2-\nu)/[N_{c}\ln(\nu^{-1}-1)]$. Rewriting Eq. (1) as $\bar{d}^{2}_{\bm{k}+\bm{q},\bm{k}}=1-|\Lambda(\bm{k},\bm{q})|^{2}$, $T_{c\bm{q}}$ with fixed band filling $\nu$ and coupling constant $g$ can be derived as
@@
in which the average of the AQD over the Brillouin zone within the band is defined as $\bar{d}^{2}_{\mathcal{A}}(\bm{q})=\sum_{\bm{k}}\bar{d}^{2}_{\bm{k}+\bm{q},\bm{k}}/N_{c}$, and $T_{0}$ is a constant which is independent of $\bm{q}$. The highest critical temperature should be determined by $T_{c\bm{Q}}=T_{0}\{1-\mathrm{min}[\bar{d}^{2}_{\mathcal{A}}(\bm{q})]\}$ over different $\bm{q}$ in the first Brillouin zone, and the corresponding optimal momentum is $\bm{q}=\bm{Q}$. We have clearly seen that the flat-band FFLO instability associated with finite-momentum $\bm{Q}$ is solely determined by the quantum geometric quantity [$\bar{d}^{2}_{\mathcal{A}}(\bm{q})$] from Eq. (6), and is independent of the interaction strength and band filling. This is the main result of our work.
In a $\mathcal{T}$-symmetric system, $u_{\bm{k}\uparrow}=u^{*}_{-\bm{k}\downarrow}$ renders $\bar{d}_{\bm{k}+\bm{q},\bm{k}}$ being the ordinary quantum distance. Since $\bar{d}^{2}_{\bm{k},\bm{k}}=0$, the quantity $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$ always has a global minimum at $\bm{q}=0$. However, as we will discuss, the presence of QGD explicitly breaks the $\mathcal{T}$ symmetry and shifts the large pair susceptibility away from $\bm{q}=0$, allowing the global minimum of $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$ to occur at $\bm{q}\neq 0$. Therefore, with Eq. (6) in hand, it becomes clear that the concept of QGD is both natural and useful for stabilizing the FFLO state in an isolated flat band.
Model Hamiltonian.—To study the Cooper pairings with QGD, now we examine a concrete model Hofmann et al. (2022, 2023); Mao and Chowdhury (2023); Thumin and Bouzerar (2024) that allows for continuous tuning of the quantum geometry. As shown in Fig. 2(a), the $\zeta$-lattice model features two orbitals (A and B) per site on a square lattice, described by the Hamiltonian $H=\sum_{\bm{k}\sigma}\hat{c}_{\bm{k}\sigma}^{\dagger}(\mathcal{H}_{\bm{k}\sigma}-\mu_{\sigma})\hat{c}_{\bm{k}\sigma}$, where

### v3

# Flat-band Fulde-Ferrell-Larkin-Ovchinnikov State from Quantum Geometric Discrepancy

@@ Abstract
We propose a new scheme for realizing Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) Cooper pairing states within flat bands, in contrast to the conventional paradigm such as the Zeeman effect. Central to our scheme is the concept of “quantum geometric discrepancy” (QGD) that measures differences in the quantum geometry of paired electrons and drives the flat-band FFLO instability. Remarkably, we find that this instability is directly related to a quantum geometric quantity known as “anomalous quantum distance”, which formally captures QGD. To model both QGD and the anomalous quantum distance, we examine a flat-band electronic Hamiltonian with tunable spin-dependent quantum metrics. Utilizing the band-projection method, we analyze the QGD-induced FFLO instability from pairing susceptibility. Furthermore, we perform mean-field numerical simulations to obtain the phase diagram of the BCS-FFLO transition, which aligns well with our analytical results. Our work demonstrates that QGD offers a general and distinctive mechanism for stabilizing the flat-band FFLO phase.
I Introduction
@@
Understanding unconventional superconductivity is a central theme in modern condensed matter physics, with important implications for both theoretical and experimental aspects Sigrist and Ueda (1991); Mineev et al. (1999); Bauer and Sigrist (2012); Scalapino (2012). Among the numerous mechanisms extending beyond the conventional Bardeen-Cooper-Schrieffer (BCS) theory of pairing, the Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) state stands out due to its characteristic finite center-of-mass momentum Cooper pairing Fulde and Ferrell (1964); Larkin and Ovchinnikov (1965); Kinnunen et al. (2018); Agterberg et al. (2020). Various mechanisms have been explored to realize exotic FFLO states in systems characterized by dispersive bands Buzdin and Kachkachi (1997); Liu (2017); Song and Koshelev (2019); Xie and Law (2023); Yuan (2023); Nakamura et al. (2024); Wan et al. (2023).
@@
In the original FFLO formulation, a spin-singlet pairing state develops finite center-of-mass momentum $\bm{Q}$ ($|\bm{Q}|\propto B/v_{F}$) under a strong Zeeman field $B$ [see Fig. 1(a)]. However, this microscopic mechanism becomes problematic in flat-band systems due to the vanishing Fermi velocity ($v_{F}\rightarrow 0$), which would lead to an unphysical divergence of the pairing momentum $\bm{Q}$. Consequently, two fundamental questions arise: First, what mechanism can stabilize the FFLO state in the flat-band limit ($v_{F}\rightarrow 0$)? Second, how is the center-of-mass momentum $\bm{Q}$ of Cooper pairs determined in flat bands? While Fermi surface nesting typically governs $\bm{Q}$ selection in dispersive band systems, this relationship becomes ambiguous in flat bands due to the ill-defined Fermi surface.
@@
Figure 1: Schematic illustrations of (a) conventional FFLO from the Zeeman effect, and (b) flat-band FFLO from QGD. In (a), an external $B$ field induces spin-split Fermi surfaces, leading to the formation of Cooper pairs with finite momentum $\bm{Q}$, where $|\bm{Q}|\propto B/v_{F}$. In (b), the finite momentum $\bm{Q}$ is stabilized when the quantum metrics of the paired electrons are different ($\mathcal{G}_{\uparrow}\neq\mathcal{G}_{\downarrow}$) within flat bands.
@@
Building upon recent studies of the quantum geometric effects in flat-band systems Peotta and Törmä (2015); Julku et al. (2016); Liang et al. (2017); Törmä et al. (2018); Hu et al. (2019); Julku et al. (2020); Xie et al. (2020); Peri et al. (2021); He et al. (2021); Verma et al. (2021); Herzog-Arbeitman et al. (2022); Törmä et al. (2022); Huhtinen et al. (2022); Kitamura et al. (2022); Hofmann et al. (2023); Jiang and Barlas (2023); Chen and Huang (2023); Törmä (2023); Tian et al. (2023); Han et al. (2024); Ticea et al. (2024); Wang and Huang (2024); Chen and Law (2024); Yu et al. (2024a, b); Hu et al. (2025), we address the aforementioned questions by linking the FFLO instability to the quantum geometry of Bloch wave functions. For concreteness, we propose a new scheme for FFLO pairing states within flat bands, driven by “quantum geometric discrepancy” (QGD). This concept refers to the differences in the quantum geometry of the wave functions of the two electrons forming a Cooper pair. Based on the band-projection formalism, we analyze the static pairing susceptibility and construct that the critical temperature for finite-momentum pairing, which leads to FFLO instability, is directly associated with the band average of a quantum geometric quantity [see Eq. (6)]—“anomalous quantum distance” (AQD) $\bar{d}_{\bm{k}^{\prime},\bm{k}}$, formally defined as
@@
Here, $u_{\bm{k}\sigma}$ represents the lattice-periodic part of the Bloch wave function at momentum $\bm{k}$ and spin $\sigma$. AQD quantifies the overlap between the wave functions of paired electrons. In the presence of time-reversal ($\mathcal{T}$) symmetry, where $u_{\bm{k}\uparrow}=u^{*}_{-\bm{k}\downarrow}$, the quantity $\bar{d}_{\bm{k},\bm{k}^{\prime}}$ becomes the well-known Hilbert–Schmidt quantum distance Shapere and Wilczek (1989); Bužek and Hillery (1996); Dodonov et al. (2000); Rhim et al. (2020), measuring the quantum mechanical distance between two Bloch states. In this scenario, the BCS (zero-momentum) pairing state is favored. In the absence of $\mathcal{T}$ symmetry, $u_{\bm{k}\uparrow}\neq u^{*}_{-\bm{k}\downarrow}$ generally makes this conclusion invalid, and Eq. (1) will be frustrated due to the QGD, enabling a finite-momentum pairing state.
As we will substantiate, even in the flat-band limit, a finite momentum-$\bm{Q}$ pairing state can be stabilized when the pairing is frustrated by QGD, specifically as $\mathcal{G}_{\uparrow}\neq\mathcal{G}_{\downarrow}$ [see Fig. 1(b)]. Here, $\mathcal{G}_{\uparrow}$ and $\mathcal{G}_{\downarrow}$ represent the quantum metrics of the Bloch states for the paired electrons. Our work establishes a comprehensive theoretical framework for understanding FFLO states in flat-band systems.
II Effective action from band projection
To investigate the FFLO instability on the flat band, we first derive the superconducting Gaussian fluctuations using the band-projection formalism Iskin (2023); Chen and Law (2024), and this analysis is rather general when considering uniform pairing. Without loss of generality, we consider the local attractive Hubbard interaction $H_{\mathrm{int}}=-U\sum_{i\alpha}\hat{c}^{\dagger}_{i\alpha\uparrow}\hat{c}^{\dagger}_{i\alpha\downarrow}\hat{c}_{i\alpha\downarrow}\hat{c}_{i\alpha\uparrow}$, where $U>0$, $\alpha$ and $i$ denote the orbital and site index.
In order to analyze the low-energy behavior of superconductivity on the relevant isolated band using field theory, we first implement the Fourier transformation ($\hat{c}_{i\alpha\sigma}=1/\sqrt{N_{c}}\sum_{\bm{k}}e^{i\bm{k}\cdot\bm{r}_{i}}\hat{c}_{\alpha,\bm{k}\sigma}$, where $N_{c}$ is the number of unit cells) and the uniform pairing approximation Törmä et al. (2018); Huhtinen et al. (2022), then we project the interaction onto the target band by $\hat{c}_{\alpha,\bm{k}\sigma}\rightarrow u_{\bm{k}\sigma}^{*}(\alpha)\hat{a}_{\bm{k}\sigma}$, where $\hat{c}_{\alpha,\bm{k}\sigma}$ and $\hat{a}_{\bm{k}\sigma}$ denotes the orbital and band-basis electron annihilation operators. We can obtain the effective attractive interaction
@@ II Effective action from band projection
in which $\Lambda(\bm{k},\bm{q})=\sum_{\alpha}u_{\bm{k}+\bm{q}\uparrow}(\alpha)u_{-\bm{k}\downarrow}(\alpha)$ is the form factor, which encodes the information of quantum geometry. Here the effective coupling constant $g\approx U/N_{\mathrm{orb}}N_{c}$, where $N_{\mathrm{orb}}$ is the number of orbitals in each unit cell. For convenience, we naturally introduce the narrow band condition such that the attractive interaction strength $g$ is large compared to the bandwidth $W$ of the isolated flat band but small compared to the band gap $\Delta_{\mathrm{gap}}$ ($W\ll g\ll\Delta_{\mathrm{gap}}$). From Eq. (2), we can arrive at an effective one-band description in the functional integral formalism Setty et al. (2023); Altland and Simons (2023), and the effective action can be expressed as
@@
Here we have used the notations $k=\left(\bm{k},\omega_{n}\right)$ and $q=\left(\bm{q},\Omega_{m}\right)$ with $\omega_{n}$ $(\Omega_{m})$ being the fermionic (bosonic) Matsubara frequencies, and $\xi_{\bm{k}\sigma}$ is the band dispersion for spin-$\sigma$ electrons. Then we adopt the standard Hubbard-Stratonovich decoupling of Eq. (3) to derive the effective action $S[\Delta_{q}]$ (for details see Supplementary Information (SI) I.A sup ) of the superconducting order parameter field $\Delta_{q}$. By retaining the second order terms of $\Delta_{q}$ in $S[\Delta_{q}]$, the Gaussian action around the normal state saddle point $\Delta_{q}=0$ reads
$S_{G}\left[\Delta_{q}\right]=S[\Delta_{q}=0]+\sum_{q}\Gamma_{q}^{-1}|\Delta_{q}|^{2},$ (4)
where the coefficient $\Gamma_{q}^{-1}=g^{-1}-\chi^{c}_{q}$ is the Cooper-pair propagator, and the static pairing susceptibility $\chi^{c}_{\bm{q}}\equiv\chi^{c}_{q=(\bm{q},0)}$ reads
@@
Here we use the notation as $\xi_{\bm{k}}=\xi_{\bm{k}\uparrow}=\xi_{-\bm{k}\downarrow}$ and $n_{\mathrm{F}}$ is the Fermi-Dirac distribution. The number equation near the critical temperature is $N=-\beta^{-1}\partial S[\Delta_{q}=0]/\partial\mu$, where $N$ is the total particle number. This yields $2\mathcal{A}\nu=\sum_{\bm{k}}\left[1-\tanh\left(\beta\xi_{\bm{k}}/2\right)\right]$, where $\mathcal{A}$ is the area of the first Brillouin zone, $\beta^{-1}=k_{B}T$, and $\nu$ is the filling factor of the target band ($0\leq\nu\leq 1$).
III Critical temperature of FFLO instability
By solving the linearized gap equation $g\chi^{c}_{\bm{q}}=1$ together with the number equation, we can derive the critical temperature $T_{c\bm{q}}$ for the instability of a static, spatially non-uniform pairing state characterized by a finite $\bm{q}$ index. Restricting the analysis to single-$\bm{q}$ states, $\Delta_{\bm{q}}\propto\delta_{\bm{q},\bm{Q}}$, the value of $\bm{Q}$ is found at the highest $T_{c\bm{q}}$ for different $\bm{q}$. If $\bm{Q}\neq 0$ ($\bm{Q}=0$), the pairing instability is associated with the FFLO (BCS) phase.
Importantly, in the exact flat-band limit, we find $\chi^{c}_{\bm{q}}=\beta\sum_{\bm{k}}|\Lambda(\bm{k},\bm{q})|^{2}(1/2-\nu)/[N_{c}\ln(\nu^{-1}-1)]$. Rewriting Eq. (1) as $\bar{d}^{2}_{\bm{k}+\bm{q},\bm{k}}=1-|\Lambda(\bm{k},\bm{q})|^{2}$, $T_{c\bm{q}}$ with fixed band filling $\nu$ and coupling constant $g$ can be derived as
@@ III Critical temperature of FFLO instability
in which the average of the AQD over the Brillouin zone within the band is defined as $\bar{d}^{2}_{\mathcal{A}}(\bm{q})=\sum_{\bm{k}}\bar{d}^{2}_{\bm{k}+\bm{q},\bm{k}}/N_{c}$, and $T_{0}=g(1/2-\nu)/[\ln(\nu^{-1}-1)]$ is a constant which is independent of $\bm{q}$. Notice that $T_{0}\propto g$, which gets maximized at $\nu\rightarrow 1/2$ and vanishes at $\nu\rightarrow 0$ and $\nu\rightarrow 1$, so does $T_{c\bm{q}}$. When we fix $T_{0}$, the highest critical temperature should be determined by $T_{c\bm{Q}}=T_{0}\{1-\mathrm{min}[\bar{d}^{2}_{\mathcal{A}}(\bm{q})]\}$ over different $\bm{q}$ in the first Brillouin zone, and the corresponding optimal momentum is $\bm{q}=\bm{Q}$, which indicates the flat-band FFLO instability is solely determined by the quantum geometric quantity $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$. This is the main result of our work. Although Eq. (6) is derived neglecting fluctuations, we argue that after considering the corrections of number equation in the presence of Gaussian fluctuations, the optimal momentum with highest critical temperature can still be determined solely by AQD, and the emergence of FFLO instability on the whole aligns with the predictions from Eq. (6), see SI I.B sup for more details.
For a $\mathcal{T}$-symmetric system, $u_{\bm{k}\uparrow}=u^{*}_{-\bm{k}\downarrow}$ renders $\bar{d}_{\bm{k}+\bm{q},\bm{k}}$ being the ordinary quantum distance. Since $\bar{d}^{2}_{\bm{k},\bm{k}}=0$, the quantity $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$ possesses a global minimum at $\bm{q}=0$. As we will see in the following sections, the introduction of QGD explicitly breaks the $\mathcal{T}$ symmetry and potentially shifts the global minimum of $\bar{d}^{2}_{\mathcal{A}}(\bm{q})$ away from $\bm{q}=0$, implying the BCS to FFLO transition. Thus, with Eq. (6) in hand, the concept of QGD is natural in stabilizing the flat-band FFLO state.
Figure 2: (a) The $\zeta$-lattice with spin-dependent long-range hoppings. $\zeta_{\sigma}$ denotes the quantum metric of spin-$\sigma$ sector. (b) Energy dispersions of the $\zeta$-lattice model in flat-band limit ($W=0$) and with finite bandwidth ($W\neq 0$). Band gap $\Delta_{\mathrm{gap}}$ is much lager than bandwidth $W$.
IV Model Hamiltonian
To exemplify the Cooper pairings with QGD, we analyze a concrete model Hofmann et al. (2022, 2023); Mao and Chowdhury (2023) that allows for continuous tuning of the quantum geometry. As shown in Fig. 2(a), the $\zeta$-lattice model features two equivalent orbitals (A and B) per site on a square lattice, described by the Hamiltonian $H=\sum_{\bm{k}\sigma}\hat{c}_{\bm{k}\sigma}^{\dagger}(\mathcal{H}_{\bm{k}\sigma}-\mu_{\sigma})\hat{c}_{\bm{k}\sigma}$, where